OpenStudy (anonymous):

p, q , r , s, n and m are positive integers such that 4p = 5q = 6r = 12s = 8n = 9m which of the following pairs contains a number , which is not an integer? {p/m , m/q} {q/9 , rs/p}

OpenStudy (anonymous):

how to find pq ...

OpenStudy (anonymous):

(p/m) is not integer \[pq=(4p ^{2})/5\]

OpenStudy (ash2326):

4p = 5q = 6r = 12s = 8n = 9m @surabhi123 we don't have to find p, q, r, s, n, m from the given data, we have \[4p=9m\] Do you get this?

OpenStudy (anonymous):

yup then

OpenStudy (ash2326):

so we get \[4p=9m\] \[\frac p m= \frac 9 4\] is p/m an integer?

OpenStudy (anonymous):

ok p/m and m/q are not integers but wat abt other ones

OpenStudy (ash2326):

we can't know q/9, unless we find q Let's check the other no. rs/p?

OpenStudy (ash2326):

4p=6r \[\frac 4 6 =\frac rp\] \[\frac {rs} p\ =\frac {4s}{6} \] now let's check if this is an integer or not are you here?